### v1

# Naoyuki Kamiyama

Abstract
The stable roommates problem is a non-bipartite version of the well-known stable matching problem. Teo and Sethuraman proved that, for each instance of the stable roommates problem in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a stable matching in the given instance. The aim of this paper is to extend the result of Teo and Sethuraman to the stable roommates problem with ties. More concretely, we prove that, for each instance of the stable roommates problem with ties in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a super-stable matching in the given instance.
1 Introduction
The stable roommates problem is a non-bipartite version of the stable matching problem in a bipartite graph [7]. In contrast to the stable matching problem, it is known that there exists an instance of the stable roommates problem where a stable matching does not exist [7, Example 3]. Thus, the problem of checking the existence of a stable matching in a given instance of the stable roommates problem is one of the most important problems in the study of the stable roommates problem. For this problem, Irving [11] proposed a polynomial-time algorithm. This algorithm is combinatorial, i.e., this does not need solving linear programs. On the other hand, in the study of the stable matching problem, linear programming approaches have been actively studied (see, e.g., [4, 13, 17]). Thus, it is natural to investigate a linear programming approach to the stable roommates problem. In this direction, Teo and Sethuraman [15] proposed the following linear programming approach. They proved that, for each instance of the stable roommates problem in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a stable matching in the given instance. Since a linear program such that the separation problem can be solved in polynomial time can be solved in polynomial time [8], the result of Teo and Sethuraman [15] can be regarded as another proof of the polynomial-time solvability of the problem of checking the existence of a stable matching in a given instance of the stable roommates problem in the complete graphs. (See [14, 16] for related work of [15] and [1, 2] for another linear programming approach to the stable roommates problem.)
In this paper, we extend the linear programming approach of Teo and Sethuraman [15] to the stable roommates problem with ties. In particular, we focus on super-stability in the stable roommates problem with ties (see Section 2 for its formal definition). Irving and Manlove [12] proposed a combinatorial polynomial-time algorithm for the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties (see also [5, 6]). In this paper, we prove that, for each instance of the stable roommates problem with ties in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a super-stable matching in the given instance. Furthermore, we prove that the separation problem for the above linear inequality system can be solved in polynomial time, and a super-stable matching can be constructed from a feasible solution to the system in polynomial time. Thus, the result of this paper gives another proof of the polynomial-time solvability of the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties in the complete graphs.
Our proof basically follows the proof of [15]. The most remarkable difference between our proof and that in [15] is the following point. In the proof of [15], we first find a feasible solution to the linear inequality system. After that, for each vertex, we arrange the incident edges having positive values in decreasing order according to its preference. In the setting of [15], since the preferences do not contain ties, this ordering for each vertex is uniquely determined. However, in our setting, since the preferences contain ties, this idea cannot be straightforwardly applied. In this paper, we resolve this issue by proving the result corresponding to the self-duality result in [2, 15] for a linear programming formulation of the stable roommates problem without ties. The proof of our result is basically the same as the self-duality result in [9] for a linear programming formulation of the super-stable matching problem in bipartite graphs. By using this self-duality result, we prove that, even if the preferences contain ties, for each vertex, the incident edges having positive values can be uniquely arranged in decreasing order according to its preference (see Lemma 3).
2 Preliminaries
Let $\mathbb{R}_{+}$ denote the set of non-negative real numbers. For each finite set $U$, each vector $x\in\mathbb{R}_{+}^{U}$, and each subset $W\subseteq U$, we define $x(W):=\sum_{u\in W}x(u)$. For each positive integer $z$, we define $[z]:=\{1,2,\dots,z\}$.
In this paper, we are given the complete graph $G=(V,E)$ on a finite vertex set $V$ such that $|V|$ is an even positive integer. Notice that $E=\{e\subseteq V\mid|e|=2\}$. For each subset $F\subseteq E$ and each vertex $v\in V$, we define $F(v)\coloneqq\{e\in F\mid v\in e\}$. For each vertex $v\in V$, we are given a transitive binary relation $\succsim_{v}$ on $E(v)$ such that, for each pair of edges $e,f\in E(v)$, at least one of $e\succsim_{v}f$, $f\succsim_{v}e$ holds. For each vertex $v\in V$ and each pair of edges $e,f\in E(v)$, if $e\succsim_{v}f$ and $f\not\succsim_{v}e$ (resp. $e\succsim_{v}f$ and $f\succsim_{v}e$), then we write $e\succ_{v}f$ (resp. $e\sim_{v}f$). Intuitively speaking, if $e\succ_{v}f$, then $v$ prefers $e$ to $f$. If $e\sim_{v}f$, then $v$ is indifferent between $e$ and $f$.
Definition 1.
A subset $\mu\subseteq E$ is called a matching in $G$ if $|\mu(v)|=1$ for every vertex $v\in V$.
For each matching $\mu$ in $G$ and each vertex $v\in V$, we do not distinguish between $\mu(v)$ and the edge in $\mu(v)$.
Definition 2.
Let $\mu$ be a matching in $G$. Then for each edge $e\in E\setminus\mu$ and each vertex $v\in e$, we say that $e$ weakly blocks $\mu$ on $v$ if $e\succsim_{v}\mu(v)$. Furthermore, for each edge $e\in E\setminus\mu$, we say that $e$ weakly blocks $\mu$ if $e$ weakly blocks $\mu$ on $v$ for every vertex $v\in e$.
Definition 3.
Let $\mu$ be a matching in $G$. Then $\mu$ is said to be super-stable if no edge in $E\setminus\mu$ weakly blocks $\mu$.
Let $v$ be a vertex in $V$, and let $e$ be an edge in $E(v)$. For each symbol $\odot\in\{\succ_{v},\succsim_{v},\sim_{v}\}$, we define $E[\mathop{\odot}e]$ (resp. $E[e\mathop{\odot}]$) as the set of edges $f\in E(v)$ such that $f\mathop{\odot}e$ (resp. $e\mathop{\odot}f$).
For a positive integer $k$, a sequence $(v_{0},v_{1},\dots,v_{k})$ of vertices in $V$ is called a walk in $G$ if $v_{i-1}\neq v_{i}$ holds for every integer $i\in[k]$. It should be noted that a walk in $G$ can pass through the same vertex more than once. For each walk $C=(v_{0},v_{1},\dots,v_{k})$ in $G$, if $v_{0}=v_{k}$, then $C$ is called a closed walk in $G$. For every closed walk $C=(v_{0},v_{1},\dots,v_{k})$ in $G$, we have $k\geq 2$.
Let $C=(v_{0},v_{1},\dots,v_{k})$ be a closed walk in $G$. For each edge $e\in E$, the multiplicity ${\sf mult}_{C}(e)$ of $e$ with respect to $C$ is defined as the number of integers $i\in[k]$ such that $\{v_{i-1},v_{i}\}=e$. The multiplicity ${\sf mult}(C)$ of $C$ is defined as $\max_{e\in E}{\sf mult}_{C}(e)$. If $k\geq 3$ and $v_{i}\neq v_{j}$ holds for every pair of distinct integers $i,j\in[k]$, then $C$ is called a cycle in $G$. Notice that, for every cycle $C$ in $G$, we have ${\sf mult}(C)=1$. If $C$ is a cycle in $G$, then we define
$V(C)\coloneqq\{v_{1},v_{2},\dots,v_{k}\},\ \ E(C)\coloneqq\{\{v_{i-1},v_{i}\}\mid i\in[k]\}.$
A set $\mathcal{C}$ of cycles in $G$ is said to be vertex-disjoint if we have $V(C)\cap V(C^{\prime})=\emptyset$ for every pair of distinct cycles $C,C^{\prime}\in\mathcal{C}$. For each subset $F\subseteq E$ and each set $\mathcal{C}$ of vertex-disjoint cycles in $G$, $\mathcal{C}$ is called a decomposition of $F$ if $\bigcup_{C\in\mathcal{C}}E(C)=F$.
For each closed walk $(v_{0},v_{1},\dots,v_{k})$ in $G$, we define $v_{k+1}\coloneqq v_{1}$ and $v_{k+2}\coloneqq v_{2}$.
Definition 4.
A closed walk $(v_{0},v_{1},\dots,v_{k})$ in $G$ is said to be dangerous if, for every integer $i\in[k]$, we have $\{v_{i+1},v_{i}\}\succsim_{v_{i}}\{v_{i},v_{i-1}\}$.
Define $\mathcal{D}$ as the set of dangerous closed walks $C$ in $G$ such that ${\sf mult}(C)\leq 4|E|$. Notice that $|\mathcal{D}|$ is finite. Then we define ${\bf P}$ as the set of vectors $x\in\mathbb{R}_{+}^{E}$ satisfying the following conditions.
$x(E(v))=1\ \ \ \mbox{($\forall v\in V$)}.$ (1)
$x(e)+\sum_{v\in e}x(E[\succ_{v}e])\geq 1\ \ \ \mbox{($\forall e\in E$)}.$ (2)
${\sum_{i\in[k]}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})\leq\lfloor k/2\rfloor}\ \ \ \mbox{($\forall(v_{0},v_{1},\dots,v_{k})\in\mathcal{D}$)},$ (3)
where we define $e_{i}\coloneqq\{v_{i-1},v_{i}\}$ for each integer $i\in[k+1]$.
The goal of this paper is to prove the following theorem.
Theorem 1.
There exists a super-stable matching in $G$ if and only if ${\bf P}\neq\emptyset$.
If the preferences do not contain ties, then Theorem 1 coincides with [15, Theorem 4]. Since we can prove that the separation problem for ${\bf P}$ can be solved in polynomial time in a similar way as [15] (see Section 4), Theorem 1 gives another proof of the polynomial-time solvability of the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties in the complete graphs. Furthermore, the proof of Theorem 1 implies that if ${\bf P}\neq\emptyset$, then a super-stable matching can be constructed from an element in ${\bf P}$ in polynomial time.
3 Proof of Theorem 1
In this section, we give the proof of Theorem 1.
For each matching $\mu$ in $G$, we define the vector $x_{\mu}\in\{0,1\}^{E}$ by
$x_{\mu}(e)\coloneqq\begin{cases}1&\mbox{if $e\in\mu$}\\ 0&\mbox{if $e\in E\setminus\mu$}.\end{cases}$
That is, $x_{\mu}$ is the characteristic vector of $\mu$.
Lemma 1.
If there exists a super-stable matching $\mu$ in $G$, then $x_{\mu}\in{\bf P}$.
Proof.
Assume that there exists a super-stable matching $\mu$ in $G$.
(1) Since $\mu$ is a matching in $G$, $x_{\mu}(E(v))=|\mu(v)|=1$ for every vertex $v\in V$.
(2) Assume that there exists an edge $f\in E$ such that (2) is not satisfied. Then (1) implies that
$1<\sum_{v\in f}x_{\mu}(E(v))-x_{\mu}(f)-\sum_{v\in f}x_{\mu}(E[\succ_{v}f])=x_{\mu}(f)+\sum_{v\in f}x_{\mu}(E[f\succsim_{v}]\setminus\{f\}).$
Thus, since $x_{\mu}\in\{0,1\}^{E}$, we have
$x_{\mu}(f)+\sum_{v\in f}x_{\mu}(E[f\succsim_{v}]\setminus\{f\})\geq 2.$ (4)
Claim 1.
$x_{\mu}(f)=0$, i.e., $f\notin\mu$.
Proof.
Assume that $x_{\mu}(f)=1$, i.e., $f\in\mu$. Then (4) implies that there exist a vertex $v\in f$ and an edge $g\in E(v)\setminus\{f\}$ such that $g\in\mu$. However, since $v\in f\cap g$ and $f\neq g$, this contradicts the fact that $\mu$ is a matching in $G$. ∎
Notice that Claim 1 and (4) imply that $f\succsim_{v}\mu(v)$ for every vertex $v\in f$. Thus, $f$ weakly blocks $\mu$. This contradicts the fact that $\mu$ is super-stable.
(3) Let $(v_{0},v_{1},\dots,v_{k})$ be a closed walk in $\mathcal{D}$. Then for each integer $i\in[k+2]$, we define $e_{i}\coloneqq\{v_{i-1},v_{i}\}$. We first prove that, for every integer $i\in[k]$,
$x_{\mu}(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})+x_{\mu}(E[e_{i+1}\succsim_{v_{i+1}}]\setminus\{e_{i+2}\})\leq 1.$ (5)
Assume that there exists an integer $j\in[k]$ such that (5) is not satisfied. Since $x_{\mu}\in\{0,1\}^{E}$,
$x_{\mu}(E[e_{j}\succsim_{v_{j}}]\setminus\{e_{j+1}\})+x_{\mu}(E[e_{j+1}\succsim_{v_{j+1}}]\setminus\{e_{j+2}\})\geq 2.$ (6)
Claim 2.
$x_{\mu}(e_{j+1})=0$, i.e., $e_{j+1}\notin\mu$.
Proof.
Assume that $x_{\mu}(e_{j+1})=1$, i.e., $e_{j+1}\in\mu$. Since $x_{\mu}(E(v))=1$ for every vertex $v\in V$,
$x_{\mu}(E[e_{j}\succsim_{v_{j}}]\setminus\{e_{j+1}\})\geq 1.$
Thus, there exists an edge $f\in E(v_{j})\setminus\{e_{j+1}\}$ such that $f\in\mu$. Since $v_{j}\in e_{j+1}\cap f$ and $f\neq e_{j+1}$, this contradicts the fact that $\mu$ is a matching in $G$. ∎
Furthermore, (6) implies that $e_{j}\succsim_{v_{j}}\mu(v_{j})$ and $e_{j+1}\succsim_{v_{j+1}}\mu(v_{j+1})$. Since the definition of a dangerous walk implies that $e_{j+1}\succsim_{v_{j}}e_{j}$, we have $e_{j+1}\succsim_{v_{j}}\mu(v_{j})$. Thus, Claim 2 implies that $e_{j+1}$ weakly blocks $\mu$. This is a contradiction. This completes the proof of (5).
It follows from (5) that
$\begin{split}k&\geq\sum_{i\in[k]}\big{(}x_{\mu}(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})+x_{\mu}(E[e_{i+1}\succsim_{v_{i+1}}]\setminus\{e_{i+2}\})\big{)}\\ &=2\sum_{i\in[k]}x_{\mu}(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\}).\end{split}$
Since $x_{\mu}\in\{0,1\}^{E}$, this implies that (3) is satisfied. This completes the proof. ∎
Lemma 1 implies that if there exists a super-stable matching in $G$, then ${\bf P}\neq\emptyset$. Thus, we prove the other direction. To this end, we first prove the following lemma. The proof of (S1) of Lemma 2 is the same as the proof of [10, Lemma 23] for bipartite graphs. Furthermore, (S2) of Lemma 2 follows from (S1) and the complementary slackness theorem of linear programming (see, e.g., [3, Theorem 3.8]). For completeness, we give the proof of Lemma 2.
Lemma 2.
For every vector $x\in\mathbb{R}_{+}^{E}$ satisfying (1), (2) and every edge $e\in E$ such that $x(e)>0$, the following statements hold.
$x(e)+\sum_{v\in e}x(E[\succ_{v}e])=1$.
For every vertex $v\in e$ and every edge $f\in E[e\sim_{v}]\setminus\{e\}$, we have $x(f)=0$.
Proof.
Consider the following linear program.
[…]
Then the dual problem of (7) is described as follows.
$\begin{array}[]{cl}\mbox{Minimize}&\alpha(V)-\beta(E)\vspace{2mm}\\ \mbox{subject to}&\displaystyle{\sum_{v\in e}\alpha(v)-\beta(e)-\sum_{v\in e}\beta(E[e\succ_{v}])\geq 1}\ \ \ \ \ \mbox{($\forall e\in E$)}\vspace{1mm}\\ &(\alpha,\beta)\in\mathbb{R}^{V}\times\mathbb{R}_{+}^{E}.\end{array}$ (8)
For each feasible solution $x$ to (7), we define $\alpha_{x}\in\mathbb{R}^{V}$ by $\alpha_{x}(v)\coloneqq x(E(v))$.
Claim 3.
For every feasible solution $x$ to (7), $(\alpha_{x},x)$ is a feasible solution to (8).
Proof.
Let $x$ be a feasible solution to (7). Then for every edge $e\in E$, since $x\in\mathbb{R}_{+}^{E}$,
$\begin{split}&\sum_{v\in e}\alpha_{x}(v)-x(e)-\sum_{v\in e}x(E[e\succ_{v}])=\sum_{v\in e}x(E(v))-x(e)-\sum_{v\in e}x(E[e\succ_{v}])\\ &=x(e)+\sum_{v\in e}x(E[\succsim_{v}e]\setminus\{e\})\geq x(e)+\sum_{v\in e}x(E[\succ_{v}e])\geq 1.\end{split}$
This completes the proof. ∎
Claim 4.
For every feasible solution $x$ to (7), $x$ and $(\alpha_{x},x)$ are optimal solutions to (7) and (8), respectively.
Proof.
For every feasible solution $x$ to (7), since
$\alpha_{x}(V)-x(E)=\sum_{v\in V}x(E(v))-x(E)=2x(E)-x(E)=x(E),$
Claim 3 and the duality theorem of linear programming (see, e.g., [3, Theorem 3.7]) imply that $x$ and $(\alpha_{x},x)$ are optimal solutions to (7) and (8), respectively. ∎
Let $x$ be a vector in $\mathbb{R}_{+}^{E}$ satisfying (1) and (2). Then Claim 4 implies that $x$ and $(\alpha_{x},x)$ are optimal solutions to (7) and (8), respectively. Let $e$ be an edge in $E$ such that $x(e)>0$. The complementary slackness theorem of linear programming (see, e.g., [3, Theorem 3.8]) implies (S1) and
$\sum_{v\in e}\alpha_{x}(v)-x(e)-\sum_{v\in e}x(E[e\succ_{v}])=x(e)+\sum_{v\in e}x(E[\succsim_{v}e]\setminus\{e\})=1.$
These statements imply that
$\sum_{v\in e}x(E[\sim_{v}e]\setminus\{e\})=0.$
Thus, since $x\in\mathbb{R}_{+}^{E}$, this implies (S2). ∎
In what follows, we fix a feasible solution $x$ to ${\bf P}$. Then we prove that we can construct a super-stable matching in $G$ from $x$. This completes the proof of Theorem 1.
For each vertex $v\in V$, we define $S(v)$ as the set of edges $e\in E(v)$ such that $x(e)>0$.
Lemma 3.
For every vertex $v\in V$ and every pair of distinct edges $e,f\in S(v)$, exactly one of $e\succ_{v}f$, $f\succ_{v}e$ holds.
Proof.
If there exist a vertex $v\in V$ and distinct edges $e,f\in S(v)$ such that $e\sim_{v}f$, then since $e,f\in S(v)$, this contradicts (S2) of Lemma 2. ∎
Lemma 4.
For each vertex $v\in V$, an edge $e\in S(v)$ satisfying
$x(E[\succ_{v}e])<1/2\leq x(E[\succsim_{v}e])$ (9)
is uniquely determined.
Proof.
Let $v$ be a vertex in $V$. Then (1) implies that there exists an edge in $S(v)$ satisfying (9). Assume that there exist distinct edges $e,f\in E(v)$ satisfying (9). Then Lemma 3 implies that exactly one of $e\succ_{v}f$, $f\succ_{v}e$ holds. Assume that $e\succ_{v}f$. Then $x(E[\succ_{v}f])\geq x(E[\succsim_{v}e])\geq 1/2$. This contradicts the fact that $f$ satisfies (9). We can prove the case where $e\succ_{v}f$ in the same way. This completes the proof. ∎
For each vertex $v\in V$, we define $d(v)$ as the edge $e\in S(v)$ satisfying (9). Lemma 4 implies that $d(v)$ is well-defined. For each vertex $v\in V$, if $d(v)=\{v,w\}$, then we define $m(v)\coloneqq w$.
Lemma 5.
For every vertex $v\in V$ such that $x(E[\succsim_{v}d(v)])>1/2$, $m(m(v))=v$.
Proof.
Let $v$ be a vertex in $V$. Define $w\coloneqq m(v)$. Then $d(v)=\{v,w\}$. Assume that $m(w)\neq v$. Thus, since $d(w)\neq d(v)$, we have at least one of $x(E[\succsim_{w}d(v)])<1/2$, $x(E[\succ_{w}d(v)])\geq 1/2$.
If $x(E[\succsim_{w}d(v)])<1/2$, then since (9) implies that $x(E[\succ_{v}d(v)])<1/2$,
$1>x(E[\succ_{v}d(v)])+x(E[\succsim_{w}d(v)])\geq x(d(v))+x(E[\succ_{v}d(v)])+x(E[\succ_{w}d(v)]).$
However, this contradicts (2). Furthermore, if $x(E[\succ_{w}d(v)])\geq 1/2$, then since the assumption of this lemma implies that $x(E[\succsim_{v}d(v)])>1/2$, we have
$1<x(E[\succsim_{v}d(v)])+x(E[\succ_{w}d(v)])=x(d(v))+x(E[\succ_{v}d(v)])+x(E[\succ_{w}d(v)]),$
where the equation follows from (S2) of Lemma 2. However, since $d(v)\in S(v)$, this contradicts (S1) of Lemma 2. ∎
Define $V^{\ast}$ as the set of vertices $v\in V$ such that $m(m(v))=v$. Then Lemma 5 implies that, for every vertex $v\in V\setminus V^{\ast}$, $x(E[\succsim_{v}d(v)])=1/2$.
Lemma 6.
For every vertex $v\in V$, if $x(E[\succsim_{v}d(v)])=1/2$, then $x(E[\succ_{m(v)}d(v)])=1/2$.
Proof.
Let $v$ be a vertex in $V$ such that $x(E[\succsim_{v}d(v)])=1/2$. Then since $d(v)\in S(v)$, it follows from (S1) and (S2) of Lemma 2 that
$x(E[\succsim_{v}d(v)])+x(E[\succ_{m(v)}d(v)])=x(d(v))+x(E[\succ_{v}d(v)])+x(E[\succ_{m(v)}d(v)])=1.$
Thus, since $x(E[\succsim_{v}d(v)])=1/2$, we have $x(E[\succ_{m(v)}d(v)])=1/2$. ∎
Lemma 7.
For every vertex $v\in V\setminus V^{\ast}$, $m(v)\in V\setminus V^{\ast}$.
Proof.
Let $v$ be a vertex in $V\setminus V^{\ast}$. Assume that $m(v)\in V^{\ast}$. Define $w\coloneqq m(v)$ and $u\coloneqq m(w)$. Then since $w\in V^{\ast}$, $m(u)=w$. Define $e\coloneqq\{v,w\}$ and $f\coloneqq\{w,u\}$.
Lemma 6 implies that $x(E[\succ_{w}e])=1/2$. Since $f=d(w)$, $x(E[\succsim_{w}f])\geq 1/2$. Assume that $x(E[\succsim_{w}f])>1/2$. If $f\succ_{w}e$, then
1/$21/2<x(E[\succsim_{w}f])\leq x(E[\succ_{w}e])=1/$ 2 .
This is a contradiction. Thus, $e\succsim_{w}f$. This implies that $x(E[\succ_{w}f])\geq x(E[\succ_{w}e])=1/2$. This contradicts the fact that $f=d(w)$. Thus, $x(E[\succsim_{w}f])=1/2$.
Since $x(E[\succsim_{w}f])=1/2$, Lemma 6 implies that $x(E[\succ_{u}f])=1/2$. However, since $f=d(u)$, (9) implies that $x(E[\succ_{u}f])<1/2$. This is a contradiction. ∎
Lemma 8.
For every vertex $v\in V\setminus V^{\ast}$, there exists exactly one vertex $w\in V\setminus V^{\ast}$ such that $m(w)=v$.
Proof.
Lemma 7 implies that it suffices to prove that, for every vertex $v\in V\setminus V^{\ast}$, the number of vertices $w\in V\setminus V^{\ast}$ such that $m(w)=v$ is at most one. Let $v$ be a vertex in $V\setminus V^{\ast}$. Assume that there exist vertices $w,u\in V\setminus V^{\ast}$ such that $w\neq u$ and $m(w)=m(v)=v$.
Define $e\coloneqq\{w,v\}$ and $f\coloneqq\{u,v\}$. Lemma 6 implies that $x(E[\succ_{v}e])=x(E[\succ_{v}f])=1/2$. Since $e,f\in S(v)$, Lemma 3 implies that $e\not\sim_{v}f$. Without loss of generality, we assume that $e\succ_{v}f$. In this case, $e\notin E[\succ_{v}e]$ and $e\in E[\succ_{v}f]$. Thus, since $E[\succ_{v}e]\subseteq E[\succ_{v}f]$ follows from $e\succ_{v}f$, $E[\succ_{v}e]$ is a subset of $E[\succ_{v}f]\setminus\{e\}$. Since $x(e)>0$, this implies that
$x(E[\succ_{v}e])\leq x(E[\succ_{v}f]\setminus\{e\})<x(E[\succ_{v}f]).$
However, this contradicts the fact that $x(E[\succ_{v}e])=x(E[\succ_{v}f])$. ∎
Let $v$ be a vertex in $V\setminus V^{\ast}$. Lemma 8 implies that there exists the unique vertex $w\in V\setminus V^{\ast}$ such that $m(w)=v$. Define $p(v)\coloneqq w$ and $d^{-}(v)\coloneqq d(w)$.
For each vertex $v\in V^{\ast}$, we define $d^{-}(v)\coloneqq d(v)$.
Lemma 9.
For every vertex $v\in V\setminus V^{\ast}$, we have $x(E[\succ_{v}d^{-}(v)])=1/2$.
Proof.
Lemma 6 implies that, for every vertex $v\in V\setminus V^{\ast}$,
$x(E[\succ_{v}d^{-}(v)])=x(E[\succ_{m(p(v))}d(p(v))])=1/2.$
This completes the proof. ∎
Lemma 10.
For every vertex $v\in V\setminus V^{\ast}$, we have $d(v)\succ_{v}d^{-}(v)$.
Proof.
Let $v$ be a vertex in $V\setminus V^{\ast}$. Lemma 9 implies that $x(E[\succ_{v}d^{-}(v)])=1/2$. If $d^{-}(v)\succsim_{v}d(v)$, then since $x(d(v))>0$ follows from $d(v)\in S(v)$, we have
$x(E[\succsim_{v}d(v)])\geq x(E[\succ_{v}d(v)])+x(d(v))>x(E[\succ_{v}d(v)])\geq x(E[\succ_{v}d^{-}(v)])=1/2.$
However, Lemma 5 implies that $x(E[\succsim_{v}d(v)])=1/2$. This is a contradiction. ∎
Define the subset $L\subseteq E$ as follows. For each pair of distinct vertices $v,w\in V$, $\{v,w\}\in L$ if and only if $m(v)=w$. (Notice that even if $m(v)=w$ and $m(w)=v$ for some pair of distinct vertices $v,w\in V$, $L$ contains a single edge between $v$ and $w$.) Then $\bigcup_{e\in L}e=V$. By considering the directed graph such that its vertex set is $V$ and its arc set contains an arc from $v$ to $m(v)$ for each vertex $v\in V$, we can prove that there exist a set $\mathcal{C}$ of vertex-disjoint cycles in $G$ and a set $M$ of pairwise disjoint edges in $L$ satisfying the following conditions.
  • $\bigcup_{e\in M}e=V^{\ast}$.
  • $\bigcup_{C\in\mathcal{C}}V(C)=V\setminus V^{\ast}$.
  • $\mathcal{C}$ is a decomposition of $L\setminus M$.
  • For every cycle $(v_{0},v_{1},\dots,v_{k})\in\mathcal{C}$ and every integer $i\in[k]$, we have $m(v_{i})=v_{i+1}$.
Lemma 11.
Let $C=(v_{0},v_{1},\dots,v_{k})$ be a cycle in $\mathcal{C}$.
For every integer $i\in[k]$, $\{v_{i},v_{i+1}\}\succ_{v_{i}}\{v_{i-1},v_{i}\}$.
$k$ is even.
Proof.
(i) For every integer $i\in[k]$, $\{v_{i-1},v_{i}\}=d^{-}(v_{i})$ and $\{v_{i},v_{i+1}\}=d(v_{i})$. This implies that (i) follows from Lemma 10.
(ii) Define $e_{i}\coloneqq\{v_{i-1},v_{i}\}$ for each integer $i\in[k+1]$. Then for every integer $i\in[k]$, since $e_{i}=d^{-}(v_{i})$,
$x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})=x(E[e_{i}\succsim_{v_{i}}])=x(E(v))-x(E[\succ_{v_{i}}e_{i}])=1-1/2=1/2,$
where the first equation follows from (i), and the third equation follows from (1) and Lemma 9. Thus, if $k$ is odd, then
$\sum_{i\in[k]}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})=k/2>\lfloor k/2\rfloor.$
However, since (i) implies that $C$ is dangerous, this contradicts (3). ∎
Since $\bigcup_{e\in L}e=V$, Lemma 11(ii) implies that we can construct a matching $\mu$ in $G$ such that $\mu\subseteq L$ by taking all the edges in $M$ and every other edge along each cycle in $\mathcal{C}$.
Lemma 12.
Let $\mu$ be a matching in $G$ such that $\mu\subseteq L$, and let $e$ be an edge in $L\setminus\mu$. Then $e$ does not weakly block $\mu$.
Proof.
There exist a cycle $(v_{0},v_{1},\dots,v_{k})\in\mathcal{C}$ and an integer $i\in[k]$ such that $e=\{v_{i-1},v_{i}\}$. In this case, $\mu(v_{i})=\{v_{i},v_{i+1}\}$. Furthermore, Lemma 11(i) implies that $\mu(v_{i})\succ_{v_{i}}e$. This implies that $e$ does not weakly block $\mu$. ∎
Lemma 13.
Let $\mu$ be a matching in $G$ such that $\mu\subseteq L$, and let $e$ be an edge in $E\setminus L$. Assume that $e$ weakly blocks $\mu(v)$. Then for every vertex $v\in e$, $d(v)\succ_{v}e\succsim_{v}d^{-}(v)$.
Proof.
We first prove the following claims.
Claim 5.
For every vertex $v\in e$, $x(E[\succ_{v}\mu(v)])\leq 1/2$.
Proof.
Let $v$ be a vertex in $e$. Since $\mu\subseteq L$, $\mu(v)\in\{d(v),d^{-}(v)\}$. If $\mu(v)=d(v)$, then (9) implies that $x(E[\succ_{v}\mu(v)])<1/2$. Assume that $\mu(v)\neq d(v)$. Then $\mu(v)=d^{-}(v)$ and $d(v)\neq d^{-}(v)$. Since $d(v)\neq d^{-}(v)$, we have $m(m(p(v)))\neq p(v)$, i.e., $p(v)\in V\setminus V^{\ast}$. This and Lemma 7 imply that $v\in V\setminus V^{\ast}$. Thus, Lemma 9 implies that $1/2=x(E[\succ_{v}d^{-}(v)])=x(E[\succ_{v}\mu(v)])$. ∎
Claim 6.
For every vertex $v\in e$, $x(E[\succ_{v}e])\leq 1/2$.
Proof.
Let $v$ be a vertex in $e$. Then since $e\succsim_{v}\mu(v)$, we have $E[\succ_{v}e]\subseteq E[\succ_{v}\mu(v)]$. This implies that $x(E[\succ_{v}e])\leq x(E[\succ_{v}\mu(v)])$. In addition, Claim 5 implies that $x(E[\succ_{v}\mu(v)])\leq 1/2$. These imply that $x(E[\succ_{v}e])\leq 1/2$. ∎
Let $v$ be a vertex in $e$. Since $\mu\subseteq L$, $\mu(v)\in\{d(v),d^{-}(v)\}$. If $v\in V^{\ast}$, then $d(v)=d^{-}(v)$. If $v\in V\setminus V^{\ast}$, then Lemma 10 implies that $d(v)\succ_{v}d^{-}(v)$. Thus, since $e\succsim_{v}\mu(v)$, $e\succsim_{v}d^{-}(v)$.
Assume that $e\succsim_{v}d(v)$. If $e\succ_{v}d(v)$, then since $x(E[\succsim_{v}e])\leq x(E[\succ_{v}d(v)])<1/2$, Claim 6 implies that
$x(e)+x(E[\succ_{v}e])+x(E[\succ_{w}e])\leq x(E[\succsim_{v}e])+x(E[\succ_{w}e])<1.$
However, this contradicts (2).
Assume that $e\sim_{v}d(v)$. Then since $e\in E\setminus L$ and $d(v)\in L$, we have $e\neq d(v)$. Thus, since $d(v)\in S(v)$, it follows from (S2) of Lemma 2 that $x(e)=0$. Furthermore, since $e\sim_{v}d(v)$, we have $E[\succ_{v}e]=E[\succ_{v}d(v)]$. Thus, since $x(E[\succ_{v}d(v)])<1/2$, Claim 6 implies that
$x(e)+x(E[\succ_{v}e])+x(E[\succ_{w}e])=x(E[\succ_{v}d(v)])+x(E[\succ_{w}e])<1.$
However, this contradicts (2). ∎
Define $B$ as the set of edges $e\in E\setminus L$ such that, for every vertex $v\in e$, $d(v)\succ_{v}e\succsim_{v}d^{-}(v)$. Then Lemmas 12 and 13 imply that, for every matching $\mu$ in $G$ such that $\mu\subseteq L$, if there exists an edge $e\in E\setminus\mu$ that weakly blocks $\mu$, then $e\in B$.
Lemma 14.
For every edge $e\in B$ and every edge $f\in M$, $e\cap f=\emptyset$.
Proof.
Assume that there exist edges $e\in B$ and $f\in M$ such that $e\cap f\neq\emptyset$. Let $v$ be a vertex in $e\cap f$. Since $f\in M$, $d(v)=d^{-}(v)$. However, this contradicts the definition of $B$. ∎
Let $C=(v_{0},v_{1},\dots,v_{k})$ be a cycle in $G$. Then for each integer $i\in[k]$ and each closed walk $C^{\prime}=(w_{0},w_{1},\dots,w_{\ell})$ in $G$ such that $w_{1}=v_{i}$, we define the closed walk $C^{+}$ in $G$ obtained from $C$ by expanding $v_{i}$ with $C^{\prime}$ by
$C^{+}\coloneqq(v_{0},v_{1},\dots,v_{i-1},w_{1},w_{2},\dots,w_{\ell},v_{i},v_{i+1},\dots,v_{k}).$
Furthermore, for each integer $i\in[k]$ and each walk $C^{\prime}=(w_{1},w_{2},\dots,w_{\ell})$ in $G$ such that $w_{1}=v_{i}$ and $w_{\ell}=v_{i+1}$, we define the closed walk $C^{+}$ in $G$ obtained from $C$ by replacing $\{v_{i},v_{i+1}\}$ with $C^{\prime}$ by
$C^{+}\coloneqq(v_{0},v_{1},\dots,v_{i-1},v_{i},w_{2},w_{3},\dots,w_{\ell-1},v_{i+1},v_{i+2},\dots,v_{k}).$
Lemma 15.
Let $C=(v_{0},v_{1},\dots,v_{k})$ be a cycle in $G$ such that $E(C)\subseteq L\cup B$. Then $k$ is even.
Proof.
Assume that $k$ is odd, and we derive a contradiction. Since $M$ is a set of pairwise disjoint edges in $L$, Lemma 14 implies that $V(C)\cap(\bigcup_{e\in M}e)=\emptyset$. First, we construct the closed walk $C^{\ast}$ in $G$ by using Algorithm 1.
1 Set $C^{\ast}_{0}\coloneqq C$.
2 for $i=1,2,\dots,k$ do
3       if $v_{i}=p(v_{i-1})$ then
4             Let $C_{i}=(w_{i,0},w_{i,1},\dots,w_{i,\ell_{i}})$ be the cycle in $\mathcal{C}$ such that $w_{i,1}=v_{i-1}$ and $w_{i,\ell_{i}}=v_{i}$. Define the walk $C_{i}^{\prime}$ in $G$ as $(w_{i,1},w_{i,2},\dots,w_{i,\ell_{i}})$.
5             Define $C^{\ast}_{i}$ as the closed walk in $G$ obtained from $C^{\ast}_{i-1}$ by replacing $\{v_{i-1},v_{i}\}$ with $C_{i}^{\prime}$.
6       else if $\{v_{i-1},v_{i}\}\in B$ then
7             Let $C_{i}=(w_{i,0},w_{i,1},,\dots,w_{i,\ell_{i}})$ be the cycle in $\mathcal{C}$ such that $w_{i,1}=v_{i}$.
8             Define $C^{\ast}_{i}$ as the closed walk in $G$ obtained from $C^{\ast}_{i-1}$ by expanding $v_{i}$ with $C_{i}$.
9       else
10             Define $C^{\ast}_{i}\coloneqq C^{\ast}_{i-1}$.
11       end if
12
13 end for
Output $C^{\ast}_{k}$ as $C^{\ast}$, and halt.
Algorithm 1 Algorithm for constructing a dangerous walk
Assume that $C^{\ast}=(c_{0},c_{1},c_{2},\dots,c_{q})$. For each integer $i\in[q+1]$, we define $g_{i}\coloneqq\{c_{i-1},c_{i}\}$. Then the following statements hold.
  • Steps 6 to 8 of Algorithm 1 imply that, for every integer $i\in[q]$, $\{g_{i},g_{i+1}\}\cap L\neq\emptyset$.
  • For every integer $i\in[q]$, if $g_{i}\in L$ (resp. $g_{i+1}\in L$), then $g_{i}=d^{-}(c_{i})$ (resp. $g_{i+1}=d(c_{i})$).
  • For every integer $i\in[q]$, $c_{i}\in V\setminus V^{\ast}$.
Claim 7.
$q$ is odd.
Proof.
Lemma 11(ii) implies that $\ell_{i}$ is even for every integer $i\in[k]$. Thus, since $k$ is odd, this claim holds. ∎
Claim 8.
For every integer $i\in[q]$, $g_{i}\succsim_{c_{i}}d^{-}(c_{i})$ and $g_{i+1}\succsim_{c_{i}}d^{-}(c_{i})$
Proof.
Let $i$ be an integer in $[q]$. If $g_{i}\in L$, then $g_{i}=d^{-}(c_{i})$. If $g_{i}\in B$, then since $c_{i}\in g_{i}$, the definition of $B$ implies that $g_{i}\succsim_{c_{i}}d^{-}(c_{i})$. If $g_{i+1}\in L$, then $g_{i+1}=d(c_{i})$. Thus, since Lemma 10 implies that $d(c_{i})\succ_{c_{i}}d^{-}(c_{i})$, $g_{i+1}\succ_{c_{i}}d^{-}(c_{i})$. If $g_{i+1}\in B$, then since $c_{i}\in g_{i+1}$, the definition of $B$ implies that $g_{i+1}\succsim_{c_{i}}d^{-}(c_{i})$. ∎
Claim 9.
For every integer $i\in[q]$, if $g_{i+1}\in B$ and $g_{i+1}\sim_{c_{i}}d^{-}(c_{i})$, then $x(g_{i+1})=0$.
Proof.
Let $i$ be an integer in $[q]$. Since $d^{-}(c_{i})\in L$ and $g_{i+1}\notin L$, we have $d^{-}(c_{i})\neq g_{i+1}$. Since $d^{-}(c_{i})\in S(c_{i})$, (S2) of Lemma 2 implies that $x(g_{i+1})=0$. This completes the proof. ∎
Claim 10.
For every integer $i\in[q]$, $E[d^{-}(c_{i})\succsim_{c_{i}}]\setminus\{g_{i+1}\}\subseteq E[g_{i}\succsim_{c_{i}}]\setminus\{g_{i+1}\}$.
Proof.
For every integer $i\in[q]$, Claim 8 implies that $E[d^{-}(c_{i})\succsim_{c_{i}}]\subseteq E[g_{i}\succsim_{c_{i}}]$. ∎
Claim 11.
For every integer $i\in[q]$, $x(E[d^{-}(c_{i})\succsim_{c_{i}}])=x(E[d^{-}(c_{i})\succsim_{c_{i}}]\setminus\{g_{i+1}\})$.
Proof.
Let $i$ be an integer in $[q]$. Then Claim 8 implies that $g_{i+1}\succsim_{c_{i}}d^{-}(c_{i})$. If $g_{i+1}\succ_{c_{i}}d^{-}(c_{i})$, then $g_{i+1}\notin E[d^{-}(c_{i})\succsim_{c_{i}}]$. Thus, we assume that $g_{i+1}\sim_{c_{i}}d^{-}(c_{i})$. In this case, if $g_{i+1}\in L$, then Lemma 10 implies that $g_{i+1}=d(c_{i})\succ_{c_{i}}d^{-}(c_{i})$. This is a contradiction. Thus, $g_{i+1}\notin L$, i.e., $g_{i+1}\in B$. In this case, Claim 9 implies that $x(g_{i+1})=0$. This completes the proof. ∎
Claim 12.
For every integer $i\in[q]$, $x(E[g_{i}\succsim_{c_{i}}]\setminus\{g_{i+1}\})\geq 1/2$.
Proof.
Let $i$ be an integer in $[q]$. Recall that $c_{i}\in V\setminus V^{\ast}$. Thus,
$\begin{split}x(E[g_{i}\succsim_{c_{i}}]\setminus\{g_{i+1}\})&\geq x(E[d^{-}(c_{i})\succsim_{c_{i}}]\setminus\{g_{i+1}\})\\ &=x(E[d^{-}(c_{i})\succsim_{c_{i}}])\\ &=1-x(E[\succ_{c_{i}}d^{-}(c_{i})])=1/2,\end{split}$
where the inequality follows from Claim 10, the first equation follows from Claim 11, the second equation follows from (1), and the third equation follows from Lemma 9. ∎
Claim 13.
$C^{\ast}\in\mathcal{D}$.
Proof.
Let $i$ be an integer in $[q]$. Recall that $\{g_{i},g_{i+1}\}\cap L\neq\emptyset$. If $g_{i},g_{i+1}\in L$, then $g_{i}=d^{-}(c_{i})$ and $g_{i+1}=d(c_{i})$. Thus, in this case, Lemma 10 implies that $g_{i+1}\succ_{c_{i}}g_{i}$. If $g_{i}\in L$ and $g_{i+1}\in B$, then since $g_{i}=d^{-}(c_{i})$, the definition of $B$ implies that $g_{i+1}\succsim_{c_{i}}g_{i}$. Assume that $g_{i}\in B$ and $g_{i+1}\in L$. Then $g_{i+1}=d(c_{i})$. Thus, the definition of $B$ implies that $g_{i+1}\succ_{c_{i}}g_{i}$.
Since $C$ is a cycle in $G$, $k\leq|V|$. Thus, since $C_{i}^{\prime}$ is a cycle in $G$ for every integer $i\in[k]$, we have ${\sf mult}(C^{\ast})\leq 2|V|\leq 4|E|$. This implies that $C^{\ast}\in\mathcal{D}$. This completes the proof. ∎
Claims 7 and 12 imply that
$\sum_{i\in[q]}x(E[g_{i}\succsim_{c_{i}}]\setminus\{g_{i+1}\})\geq q/2>\lfloor q/2\rfloor.$
Claim 13 implies that this contradicts (3). This completes the proof. ∎
We now ready to prove Theorem 1. Define the subgraph $H$ of $G$ by $H\coloneqq(V,L\cup B)$. Then Lemma 15 implies that $H$ is bipartite. Thus, there exists a partition $\{P,Q\}$ of $V$ such that, for every edge $e\in L\cup B$, $|e\cap P|=|e\cap Q|=1$.
Lemma 16.
Let $(v_{0},v_{1},\dots,v_{k})$ be a cycle in $\mathcal{C}$. Then one of the following statements holds.
  • $\{v_{1},v_{3},\dots,v_{k-1}\}\subseteq P$ and $\{v_{2},v_{4},\dots,v_{k}\}\subseteq Q$.
  • $\{v_{1},v_{3},\dots,v_{k-1}\}\subseteq Q$ and $\{v_{2},v_{4},\dots,v_{k}\}\subseteq P$.
Proof.
This lemma follows from the fact that $\{v_{i-1},v_{i}\}\in L$ for every integer $i\in[k]$. ∎
Define $\mu\coloneqq\{d(v)\mid v\in P\}$. Then Lemma 16 implies that $\mu$ is a matching in $G$. Thus, what remains is to prove that $\mu$ is super-stable. For every edge $e\in B$, since $|e\cap P|=1$, there exists a vertex $v\in e$ such that $\mu(v)=d(v)$. Thus, since the definition of $B$ implies that $d(v)\succ_{v}e$, $\mu$ is super-stable. This completes the proof of Theorem 1.
4 Polynomial-Time Solvability
Here we prove that the separation problem for ${\bf P}$ can be solved in polynomial time. The proof follows the proof in [15] for preferences without ties. The remarkable difference is that we need (S2) of Lemma 2 to prove the non-negativity of the cost function on an auxiliary directed graph (see Lemma 17).
Let $x$ be a vector in $\mathbb{R}_{+}^{E}$. It is not difficult to see that we can determine whether $x$ satisfies (1) and (2) in polynomial time. Assume that $x$ satisfies (1) and (2). In what follows, under this assumption, we consider the problem of determining whether $x$ satisfies (3).
Define the auxiliary directed graph $D=(N,K)$ as follows. Define the vertex set $N$ of $D$ by
$N\coloneqq\{\langle v,w,s\rangle,\langle w,v,s\rangle\mid\{v,w\}\in E,s\in V\setminus\{v,w\}\}\cup\{\langle v,w,v\rangle,\langle w,v,w\rangle\mid\{v,w\}\in E\}.$
Then for each pair of distinct vertices $\langle v,w,s\rangle,\langle p,q,r\rangle\in N$, there exists an arc $(\langle v,w,s\rangle,\langle p,q,r\rangle)$ from $\langle v,w,s\rangle$ to $\langle p,q,r\rangle$ in $K$ if and only if $w=p$, $s=q$, and $\{p,q\}\succsim_{w}\{v,w\}$. For each arc $(\langle v,w,s\rangle,\langle w,s,r\rangle)\in K$, we define the cost ${\sf cost}(\langle v,w,s\rangle,\langle w,s,r\rangle)$ by
${\sf cost}(\langle v,w,s\rangle,\langle w,s,r\rangle)\coloneqq 1-x(E[\{v,w\}\succsim_{w}]\setminus\{w,s\})-x(E[\{w,s\}\succsim_{s}]\setminus\{s,r\}).$
Lemma 17.
Let $(\langle v,w,s\rangle,\langle w,s,r\rangle)$ be an arc in $K$. Then ${\sf cost}(\langle v,w,s\rangle,\langle w,s,r\rangle)\geq 0$.
Proof.
Define $e\coloneqq\{v,w\}$, $f\coloneqq\{w,s\}$, and $g\coloneqq\{s,r\}$. Recall that $x$ satisfies (1) and (2). Thus, (2) for $f$ implies that
$\begin{split}0&\leq x(f)+x(E[\succ_{w}f])+x(E[\succ_{s}f])-1\\ &=x(f)+(1-x(E[f\succsim_{w}]))+(1-x(E[f\succsim_{s}]))-1\\ &=1-x(E[f\succsim_{w}])-x(E[f\succsim_{s}])+x(f),\end{split}$ (10)
where the first equation follows from (1). If $x(f)=0$, then (10) implies that
$\begin{split}0&\leq 1-x(E[f\succsim_{w}])-x(E[f\succsim_{s}])\\ &\leq 1-x(E[e\succsim_{w}])-x(E[f\succsim_{s}])\\ &\leq 1-x(E[e\succsim_{w}]\setminus\{f\})-x(E[f\succsim_{s}]\setminus\{g\})={\sf cost}(\langle v,w,s\rangle,\langle w,s,r\rangle),\end{split}$
where the second inequality follows from $f\succsim_{w}e$. Thus, we assume that $x(f)>0$. In this case, (S2) of Lemma 2 implies that $x(E[f\sim_{w}]\setminus\{f\})=x(E[f\sim_{s}]\setminus\{f\})=0$. Thus, (10) implies that
$\begin{split}0&\leq 1-x(E[f\succsim_{w}])-x(E[f\succsim_{s}])+x(f)\\ &=1-x(E[f\succ_{w}])-x(f)-x(E[f\succ_{s}])-x(f)+x(f)\\ &=1-x(E[f\succ_{w}])-x(E[f\succ_{s}])-x(f).\end{split}$ (11)
If $f\succ_{w}e$, then
$\begin{split}{\sf cost}(\langle v,w,s\rangle,\langle w,s,r\rangle)&=1-x(E[e\succsim_{w}]\setminus\{f\})-x(E[f\succsim_{s}]\setminus\{g\})\\ &\geq 1-x(E[e\succsim_{w}])-x(E[f\succsim_{s}])\\ &\geq 1-x(E[f\succ_{w}])-x(E[f\succ_{s}])-x(f).\end{split}$
If $f\sim_{w}e$, then (S2) of Lemma 2 implies that $x(E[e\sim_{w}]\setminus\{f\})=0$. Thus, since $f\sim_{w}e$,
$\begin{split}{\sf cost}(\langle v,w,s\rangle,\langle w,s,r\rangle)&=1-x(E[e\succsim_{w}]\setminus\{f\})-x(E[f\succsim_{s}]\setminus\{g\})\\ &=1-x(E[e\succ_{w}])-x(E[f\succsim_{s}]\setminus\{g\})\\ &\geq 1-x(E[e\succ_{w}])-x(E[f\succsim_{s}])\\ &=1-x(E[e\succ_{w}])-x(E[f\succ_{s}])-x(f)\\ &=1-x(E[f\succ_{w}])-x(E[f\succ_{s}])-x(f).\end{split}$
Thus, in both cases, (11) implies that ${\sf cost}(\langle v,w,s\rangle,\langle w,s,r\rangle)\geq 0$. This completes the proof. ∎
For a positive integer $k$, a sequence $(a_{0},a_{1},\dots,a_{k})$ of vertices in $N$ is called a directed walk in $D$ if we have $a_{i-1}\neq a_{i}$ and $(a_{i-1},a_{i})\in K$ for every integer $i\in[k]$. Let $W=(a_{0},a_{1},\dots,a_{k})$ be a directed walk in $D$. If $a_{0}=a_{k}$, then $W$ is called a closed directed walk in $D$. Furthermore, if $W$ is a closed directed walk in $D$ and $a_{i}\neq a_{j}$ holds for every pair of distinct integers $i,j\in[k]$, then $W$ is called a directed cycle in $D$. If $k$ is odd, then $W$ is called an odd closed directed walk in $D$ or an odd directed cycle in $D$. If $W$ is a closed directed walk in $D$, then we define the cost ${\sf cost}(W)$ of $W$ by ${\sf cost}(W)\coloneqq\sum_{i=1}^{k}{\sf cost}(a_{i-1},a_{i})$.
Lemma 18.
$x$ satisfies (3) if and only if ${\sf cost}(W)\geq 1$ for every odd directed cycle $W$ in $D$.
Proof.
Assume that there exists an odd directed cycle $W$ in $D$ such that ${\sf cost}(W)<1$, and we prove that $x$ does not satisfy (3). Assume that $W=(a_{0},a_{1},\dots,a_{k})$ and $a_{i}=\langle v_{i-1},v_{i},v_{i+1}\rangle$ for each integer $i\in[k]$. Notice that, since $(a_{i-1},a_{i})\in K$ for every integer $i\in[k]$, $v_{1},v_{2},\dots,v_{k}$ are well-defined and we have $v_{0}=v_{k},v_{k+1}=v_{1}$. Define $C\coloneqq(v_{0},v_{1},\dots,v_{k})$ and $e_{i}\coloneqq\{v_{i-1},v_{i}\}$ for each integer $i\in[k+2]$.
Claim 14.
${\sf mult}(C)\leq 2|V|\leq 4|E|$.
Proof.
Since the number of vertices $\langle v,w,s\rangle\in N$ such that $\{v,w\}=e$ is at most $2|V|$ for every edge $e\in E$, this claim follows from the fact that $W$ is a directed cycle in $D$. ∎
Since $e_{i}\succsim_{v_{i}}e_{i-1}$ follows from $(a_{i-1},a_{i})\in K$ for every integer $i\in[k]$, Claim 14 implies that $C\in\mathcal{D}$. Furthermore,
$\begin{split}1>{\sf cost}(W)&=\sum_{i=1}^{k}\big{(}1-x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})-x(E[e_{i+1}\succsim_{v_{i+1}}]\setminus\{e_{i+2}\})\big{)}\\ &=k-2\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\}).\end{split}$
Thus, since $k$ is odd,
$\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})>(k-1)/2=\lfloor k/2\rfloor.$
This implies that $x$ does not satisfy (3).
Assume that $x$ does not satisfy (3), i.e., there exists a closed walk $C=(v_{0},v_{1},\dots,v_{k})\in\mathcal{D}$ such that
$\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})>\lfloor k/2\rfloor,$ (12)
where we define $e_{i}\coloneqq\{v_{i-1},v_{i}\}$ for each integer $i\in[k+1]$. For each integer $i\in[k]$, we define $a_{i}\coloneqq\langle v_{i-1},v_{i},v_{i+1}\rangle$. Define $a_{0}\coloneqq a_{k}$ and $W\coloneqq(a_{0},a_{1},\dots,a_{k})$. Since $\{v_{i},v_{i+1}\}\succsim_{v_{i}}\{v_{i-1},v_{i}\}$ for every integer $i\in[k]$, $W$ is a closed directed walk in $D$. Define $e_{k+2}\coloneqq\{v_{1},v_{2}\}$. Lemma 17 implies that
$\begin{split}0&\leq\sum_{i\in[k]}{\sf cost}(\langle v_{i-1},v_{i},v_{i+1}\rangle,\langle v_{i},v_{i+1},v_{i+2}\rangle)\\ &=\sum_{i\in[k]}\big{(}1-x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})-x(E[e_{i+1}\succsim_{v_{i+1}}]\setminus\{e_{i+2}\})\big{)}\\ &=k-2\sum_{i\in[k]}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\}),\end{split}$ (13)
If $k$ is even, then $k/2=\lfloor k/2\rfloor$. Thus, (12) and (13) imply that $k$ is odd. This implies that $W$ is an odd closed directed walk in $D$. Furthermore,
$\begin{split}{\sf cost}(W)&=k-2\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})\\ &<k-2\lfloor k/2\rfloor=k-2((k-1)/2)=1.\end{split}$
Thus, by using $W$ and Lemma 17, we can prove that there exists an odd directed cycle $W^{\prime}$ in $D$ such that ${\sf cost}(W^{\prime})<1$ as follows. Assume that $W$ is not a directed cycle in $D$. Let $i$ be the minimum integer in $[k]$ such that there exists an integer $j\in[i-1]$ satisfying the condition that $a_{i}=a_{j}$. The definition of a directed walk implies that $j\neq i-1$. Then $W^{\circ}=(a_{j},a_{j+1},\dots,a_{i})$ is a directed cycle in $D$. Since ${\sf cost}(W)<1$, Lemma 17 implies that ${\sf cost}(W^{\circ})<1$. If $i-j$ is odd, then the proof is done. Assume that $i-j$ is even. Define the closed directed walk $W^{\bullet}$ in $D$ by
$W^{\bullet}\coloneqq(a_{0},a_{1},\dots,a_{j-1},a_{i},a_{i+1},\dots,a_{k}).$
Then Lemma 17 implies that ${\sf cost}(W^{\bullet})<1$. Since $i-j$ is even and $W^{\bullet}$ is obtained from $W$ by removing $i-j$ vertices, $W^{\bullet}$ consists of an odd number of vertices. Thus, by repeating this, we can obtain a desired odd directed cycle in $D$. This completes the proof. ∎
Lemma 18 implies that if we can find an odd directed cycle $W$ in $D$ that minimizes ${\sf cost}(W)$ in polynomial time, then the separation problem for ${\bf P}$ can be solved in polynomial time. Since it is known that Lemma 17 implies that this problem can be solved in polynomial time (see [8, Problem 8.3.6]), the separation problem for ${\bf P}$ can be solved in polynomial time.
Finally, we consider the problem of finding a super-stable matching in $G$ if one exists. First, we determine whether ${\bf P}\neq\emptyset$. If ${\bf P}=\emptyset$, then there does not exist a super-stable matching in $G$. Assume that ${\bf P}\neq\emptyset$. Then we compute an element $x\in{\bf P}$ in polynomial time [8]. For each vertex $v\in V$, we can compute $d(v)$ in polynomial time by using $x$. Furthermore, we can compute $L,B$ in polynomial time by using $d(\cdot)$. We can compute a partition $P,Q$ by the breadth-first search in polynomial time. These imply that we can find a super-stable matching in $G$ in polynomial time.